OpenStudy (anonymous):

If you pick a shirt at random from the given batch of 165 shirts, what is the probability that it is red and the size is medium? Shirt Color Size Large Medium Total Red 42 48 90 Blue 35 40 75 Total 77 88 165

OpenStudy (just_one_last_goodbye):

Can you organize it? the numbers and stuff are like right ontop of each other

OpenStudy (just_one_last_goodbye):

Color Large Medium Total(shirts) Red 42 48 90 Blue 35 40 75 Total(sizes) 77 + 88 = 165

OpenStudy (just_one_last_goodbye):

Is that what it looks like?

OpenStudy (amoodarya):

 \[\frac{48}{42+48+35+40}=p\]